Weekly Report for Week Ending April 8, 2004


 Exec. Comm. Agenda
Highlights
LSC
Administration
Hanford Observatory
Livingston Observatory
MIT
Caltech
Detector
40 Meter
TNI
LASTI
Data Analysis
Adv. LIGO Development
Past Weekly Reports

The LIGO Executive Committee Agenda for Monday  April 12, 2004 will be:

 (Meeting time: 10:30 am Pacific Time)
 

  1. Announcements
  2. LSC Issues (Saulson)
  3. Comments on Weekly Report
  4. WBS 1 LIGO I Construction (Lindquist)
  5. WBS 2 LIGO Lab Operations
  6. WBS 3 and 4  Advanced R&D and LIGO II (Shoemaker)
  7. CHANGE CONTROL BOARD/TECHNICAL REVIEW BOARD SESSION AS NEEDED


Special Items:


Special Announcements:


To LIGO:

Thank you all for a remarkable ten years building LIGO. We can all be proud of what has been accomplished. I have learned a great deal that I will take with me to my new project. But mostly I will remember you all.

Best wishes,

Gary Sanders


Weekly Report Highlights
 


LSC Issues (Saulson)


no report


LIGO Laboratory Administration (Lindquist)


Status of LSC/MOU Research Updates and Program Reports (Petrac)(LSC Research Updates through Aug. 2004,and Prog. Report through Feb. 2004)

Carleton Univ.:
Univ. of Oregon: Goddard:
LaTech:
Univ. of Texas at B.:
Moscow State:
Trinity Univ.:
Nat. Observatory of Japan-TAMA:
Univ. of Washington:
Univ. of Michigan:
Univ. of Syracuse:
Southeastern Louisiana Univ.:

LIGO Weekly Site Telecon (Lindquist)
There was a site teleconference Thursday, April 8, 2004.  The following issues were among those discussed:

The list of current actions revised to reflect the status of open actions assigned through the last update (March 4, 2004) may be found at ACTION LIST.

PROPERTY MANAGEMENT (Chargois)

From: Ed Chargois <chargois_e@ligo.caltech.edu>


DOCUMENT CONTROL CENTER (Turner, Mak)

>From: Linda Turner - turner@ligo.caltech.edu>

Web pages for the DCC give simple how-to's for document numbering, easy access to the latest on-line documents, and search capabilities for the DCC database. Take a look. . .

ACCOMPLISHMENTS

> From: Cleveland Mak <mak_c@ligo.caltech.edu>

ACTIVITY
04/08/04 Packages Faxes
In 21 20
Out 13 11

Press here to access the DOCUMENT CONTROL CENTER WEB PAGE.


COST SCHEDULE CONTROL SYSTEMS (Cunningham, Brambila, Kaufman, Salone)

From: Esther Cunningham <esther@ligo.caltech.edu>

Press here for ACCOUNTS PAYABLE HISTORY DATA.

From: "Brambila, Ruth" <Ruth.Brambila@caltech.edu>

From: Gina Salone <gsalone@ligo.caltech.edu> From: Florence Kaufman <fkaufman@ligo.caltech.edu> Financial reports can be found at: http://docuserv.ligo.caltech.edu/~fireport.
SUBCONTRACTS MANAGEMENT (Petrac, Jasnow)

From: irena@ligo.caltech.edu (Irena Petrac)

From: Ed Jasnow <jasnow@ligo.caltech.edu>

SUPPORT (Baldon, Lloyd, Tischler)

>Irene BaldonPLEASE NOTE:  This report covers only 4 working days.

>Dorothy Lloyd >From: Ryan Tischler <rtischle@ligo.caltech.edu>

ADVANCED LIGO (Cost Schedule Control Systems) T. Frey
From: Thomas Frey <tfrey@ligo.caltech.edu>
 

For list of documents that are being used to develop Adv. LIGO Cost and Schedule, see http://www.ligo.caltech.edu/~tfrey/Cost_MTG_082002/

Working on the integration of project summary data created by Carol.  Due to Carol on Friday.

Development of the Advanced LIGO Project Controls Guidebook continues.

Project Web Site for posting schedule and progress related data continues to be updated with the latest and greatest.


Reports (Lindquist)

LIGO Video--Tom Lucas (Lucas Productions) will be visiting Caltech April 22 and 23 to interview people.  Tom is building the "plot line" for a 20-25 minute NOVA type video on LIGO under a Grant from the NSF.  I will be arranging interviews with Caltech personnel for Tom.  Tom will be traveling from Caltech to the Hanford site to participate in the "star party" currently scheduled for April 24.


Change Control/Contingency (Lindquist)
There are no open change requests.

Human Resources (Akutagawa)

From: Cindy Akutagawa <cindy@ligo.caltech.edu>


Quality/Safety (Tyler)

>From: Bill Tyler <tyler@ligo.caltech.edu>
 No report.


LIGO Hanford Observatory (LHO) and Interferometer Operations (Raab)



 

Summary of Commissioning Activities at LIGO Hanford Observatory  (compiled by M. Landry)

After considerable time with one or the other machine in some marginal state, it was
good to see both IFOs posting decent ranges on the weekend (H1 at 4.4Mpc,
H2 at 1.2Mpc).  Team laser had brought the 4k PSL back online, but continued
tweaking through the week.

The current understanding of the least-significant byte problem is summarized here.  A
technical note will follow up in a week or so, detailing the arduous detective work.

4K IFO
Groundwork for the eventual installation of the output mode cleaner on
ISCT4 has begun.  Real estate is at a premium however, designs including
pathlengths as small as 6cm.  Work continues.

Dark and shot noises were measured in the 4k mode cleaner photodiode.  The
measurements are given here, plus proposed changes to the 4k mode cleaner board.
A TCS transfer function (thermal compensation system coupling to AS_Q) was
performed.  What remains to be determined is whether or not TCS laser noise will
cause trouble. A non-resonant 61.2MHz photodiode was readied and  installed on
ISCT1.

2K IFO
Oscillator phase noise coupling to AS_Q in the 2k is 30X lower than in that of
the 4k.  The AS_Q high frequency noise performance had always been best in
the 2k, but the source of this difference is not understood.

A WFS1 noise budget was worked out in part here, concentrating on low
frequencies, while at high frequencies, the wave front sensor path was shown to be
dark noise limited.
 


LIGO Livingston Observatory (LLO) and Interferometer Operations (Zucker)


---------------------------------------------------------------------
 

L1
--------------------------------------------------------------------
Major progress on the HEPI installation and moving of the electronics
into RFI-tight enclosures at the X end (phase 1 of the EMC retrofit).
Also significant advances on characterization of laser intensity noise
and regeneration of the X end cryopump.
 

LLO Outreach (B. Wooley)
---------------------------------------------------------------------
63 fifth graders from St. John School visited LIGO today for hands on
science, tour and painting.  Good time had by all.
 
 

Detector commissioning (Frolov, O'Reilly, Gretarsson)
---------------------------------------------------------------------
I worked on the ISS noise characterization. The calibration of the RIN
was corrected(was off by factor of 2) and consistency checks
between independent PDs(MC transmitted and REFL) were made. TheRIN
monitoring channel was found to be electronics noise limited <400Hz.
The low noise monitoring channel will be connected and used to monitor the
RIN. The estimates of the contribution of the intensity noise to ifo
output were updated based on improved RIN calibrations and measurements.
 

Andri, Brian, and I made measurements of the effect of the PRC servo
loop offset on the cavity gain and mode shape. Started comparing results
with Erika's FFT calculations. The caculation shows qualitative agreement
with measurement of the sideband gain(SPOB) vs prc cavity offset. We
are planning to make more measurements in both State 2 and State 4(LLO
and/or LHO) and make quantitative comparisons for cavity field
distribution, power buildup, and optical gains.
 
 

LLO Seismic retrofit (R. Abbott, O. Spjeld, K. Carter and many others)
---------------------------------------------------------------------
1.  The hydraulic valves at LLO have been leak checked by Katrina and
Tom.  The results are that we are a few short (8) of what we need at LLO
to complete.  The valves that didn't meet our test requirements have
been boxed up along with details of the failure mode, and are being
shipped back to the manufacturer.
 

2.  All the electronics is installed in the X-end station.  The last
cables are being run today.  System debug will begin immediately.
 

3.  The Y-end station is complete as far as actuator installation is
concerned.  Next week, it's on to the vertex.
 

4.  The piping installation is going well.  The plan is to finish the
BSCs first so that we stay in step with the actuator installation.  The
HAM chambers are the first to be retrofitted in the LVEA.
 

5.  The HAM lifting plate is being tested this week by Joe Lacour and
Southern Enterprises.  We will have the HAM lift plate by early (Monday)
next week.
 

6.  Actuators for the HAMs are being built by Oddvar and crew in the
staging building.
 

Oddvar adds:
---------------------------------------------------------------------
 

- performed 3D-model fit check of HAM lifting plate in SolidWorks
 

- provided Southern Enterprises with machine drawings for lifting plates
 

- revised the HEPI actuator bleeding procedure
 

- requested quote from Ameriflex for the vertex flex hoses
 

- ordered necessary hardware for HEPI installation in vertex
 

- created template for location of 4-way valve mounting plate on HAM
piers
 

- worked with Joe Lacour to prepare for the lifting test (Monday, April
12)
 

- updated assembly drawings for the HAM HEPI configuration
 

- outlined the HAM Lifting Procedure
 

- started to assemble the first HAM housings
 

Katrina adds:
---------------------------------------------------------------------
I am finally done leak testing the valves.  53 of the 55 calibrated valves
passed the leak test.  We now have six full sets of eight valves and five
spares.  The valves have been grouped into sets of eight and placed in the
clean room along the work bench (each valve has the number of the group it
belongs to in the corner of the tag).  The 29 rejected valves have been
prepared and boxed for shipment, and a leak test stand has been added to the
package also (I have ncluded a sheet with the serial numbers/problems
associated with the valves and an instruction sheet on how to use the leak
test stand).  The shipment should be picked up today for next day delivery.
 
 

HPLF and Detector (Rupal Amin)
---------------------------------------------------------------------
HPLF:
IPG has set a earliest shipping date for the 100 W laser.  We
expect shipping and deliver, at best, between April 19 and April 23. We
have also purchased a new BeamScan capable of handling the 100 W CW
laser.  The ship date for this device is April 23.  Hopefully this will
permit us to prepare for beam quality diagnostics while IPG's laser
engineer is en route.  The lab's laser interlock remains to be
finalized, and the lab needsto be cleaned prior to unpacking the 100 W.
 

2k IFO FI Installation Preparation:
According to EOT, the 2k Faraday
rotator is one day from completion.  It will be shipped either Friday or
Monday directly to UF for alignment and mount fitting. This rotator's
fields will be emprically measured to validate the qualitative effects
of the stray B-fields in the HAM 7 chamber.
 

Also UF determined that the halfwave plate (HWP) ordered from CVI was
not compatible with the UHV system.  This HWP will be returned and a new
pair of zero order HWPs from Karl-Lambrecht has been ordered.  These
plates will be ready for shipment to UF in 10 days.  However,
Karl-Lambrecht has courteously allowed us to borrow a scratched HWP for
mounting and fitting purposes.  This substandard waveplate will be
returned when the research grade HWPs reach UF.
 

PSL Work:
After last weeks reference cavity failure, I begain inspecting
several time series data sets for possible failure signal  precursors.
Finding none I have come to the default (and perhaps unsatisfactory)
conclusion that the reason for reference cavity failure Wednesday last
week was that I had "upset" the laser's control loops by turning the
laser on and off nearly 5 times in one day.  The ref. cavity has not
fallen out of lock since last week without operator intervention.
 

I have also today (Thursday) obtained a power budget of the PSL.  See
the LLO elog for details. Mode matching is to come early tomorrow if not
later today.
 

Photon Calibrator:
I have been assisting Doug in preparing the photon
calibrator.  This has mainly focused on testing the Nd:YLF laser and
mocking up the beamline.  New safety procedures will also be required to
safe the end stations.  These procedures are contained in the laser
safety plan.  Rusyl Wooley is also being informed of  the safety issues
and has designed a simple interlock that will prevent unintentional
laser discharge while the calibrator is being serviced.
 

Laser Safety Documents:
I aided Rich Riesen in the final modifications to
the LLO Laser Safety Plan. This document is now ready for review and
should contain all essential information including eye examination,
goggles, hazard areas, etc.
 
 
 

AdL optical modeling (Ken Yoshiki Franzen)
---------------------------------------------------------------------
Continued my study of the Melody AdLIGO mode cleaner model. I am
checking and testing the code written by Ray Beausoleil and Amber Lynn
Bullington as requested by the UF LIGO group. So far the model looks
rather promising even though the mirror-field mismatch and astigmatism
is not yet implemented for the curved mirror.
 
 

General Computing (Roddy)
---------------------------------------------------------------------
* Current bandwidth usage can be seen at
http://teche.ligo-la.caltech.edu/mrtg/LLO-Router/130.39.245.1_1.html
 

* Archived bandwidth usage can be seen at
http://teche.ligo-la.caltech.edu/mrtg/archive/ under the dated folder
for the week of interest.
 

* moved the mail server, file server, support web server, and another
machine to a new rack in the computer users room.  Also moved the Gb
switch into the new rack.  There is  a larger UPS in the rack also which
should run the servers for ~20-40 minutes instead of ~5 for the old one.
 

* Set up a new laptop for Bonnie.
 

* Moved all of the shared application installation files from a CD
server to our main file server.
 

* Made further firewall changes for LDAS.  This is to support the new
servers LDAS has installed at LLO.
 

* Spent a considerable amount of time installing the new rack and
rearranging the computer users room to accommodate the new rack and UPS
in the computer users room.  This has been needed for a long time.
Finally I can move all of the servers out of the Comm room and into a
proper rack.  There is one rack in the room now, later this week a
second rack will be available.  Both racks will now have better UPSs and
network switches.  I will have to order a new KVM however.
 

* Working on providing more network ports to the new building.  I have
finally run out of switch ports there.  This is causing a problem since
I need a couple more wall jacks activated.  I am working on  a short
term solution since delivery on a fiber I need is several weeks out.
 

* Setting up a couple of rack mount computers for the Intellution
software.  These will be shipped to the contractor and returned with the
necessary software installed.
 
 

Data Analysis (Yakushin)
---------------------------------------------------------------------
1) Debugging the new MDC frames script for LIGO and LIGO-TAMA analysis.
 

2) Together with S. Klimenko and M. Rakhmanov preparing GDDAW-8
presentation on waveburst for publication in Classical and Quantum
Gravity.
 

3) Helping Michele Zanolin to learn to use waveburst for the internal
burst group software review.
 
 

Vacuum Equipment (MZ for Rus Wooley, Rai Weiss and Joe Langdale)
---------------------------------------------------------------------
The second (X end) 80K cryopump was successfully regenerated. This
one evidently harbored substantially more water vapor that the Y end,
which is puzzling. After recooling the station base pressure and residual
gas spectrum are substantially better, as was found for Y. Also it again looks
like there is a significant (~ 20%) reduction in the LN2 consumption rate,
presumably due to the reduced IR emissivity of the now clean, ice-free
trap surface.
 

Plan is to reopen both ends Friday for a long
weekend of full-IFO commissioning (microseism permitting).
 

CDS support (Ash Khan)
---------------------------------------------------------------------
1) Documented the Ref Cavity Tidal servo Autotracker software.
 

2) Developed a test software to find, why the Y-End front-end code runs
slower than the X-End. The result will be posted once the X-End is up and
running.
 

3) Updated the reboot procedure of l1adcupem controller.
 


Detector/Technical Support (Coyne)



 

DETECTOR SCIENCE & ENGINEERING

Seismic Upgrade Project

see also:

Ken Mailand

Working on the task of organizing the HEPI materials and documentation, and sending drawing files to the DCC.

DMT

John Zweizig

This week I worked on documentation of DMT for DASWG. I also modified

the DMT process manager to use the new trend directory structure, and 

worked on fixing and documenting the DMT simulation package.

CDS

see also the CDS weekly meeting minutes in the commissioning archives:

CDS Software

Rolf Bork reporting

- Finishing up some HEPI work, particularly addition of ADC overflow counters. I 

will be at LLO next week for initial testing.

- ASC upgrade in progress to add MC WFS. This will be installed at LLO end of 

next week and at LHO week of April 19. 

- Received three of the four new VME Pentium4 processor boards; need to build 

new vxWorks kernel.

CDS Hardware

Ben Abbott reporting
·I have collected all of the previous suggestions and comments made regarding the RFPD redesign.  I was given a newer schematic that Rich Abbott had made back when he was doing this task.  I have incorporated some new features and have changed the schematic somewhat to include some more recent concerns like RFI mediation.  I have sent the modified schematic to Vern for comments.

Jay Heefner reporting

The EMI upgrade of the LLO x end station is underway and the controls should be operational by Friday.

Sander Liu

Working on redesigning the EO Receiver circuit board to solve the on board oscillation problem.

PSL

Peter King, Flavio Nocera

  No apparent problems have been observed with the NPRO laser that was

removed from the 40-m Lab 10-W laser.  When inspected at random intervals

the output power was within 2 mW of its setpoint and the crystal

temperature was within 0.2 degC of its setpoint.  However when the crystal

temperature and current, and the diode temperature and current signals are

displayed on an oscilloscope there is a strong ~20 MHz signal present.

This masks any glitching in the signals, if there is any so I have not been

able to compare these signals with those posted in the LLO e-log.

    A broken capacitor --- 10 pF I think --- was found on the current shunt

board.  This was replaced and made a small difference to the transfer

function around 2 MHz.

OMC

PeterKing

    The piece for the output modecleaner vacuum chamber was EDMed okay.

Part of the delay was in finding a local wire cutting shop that could

handle pieces larger than 6 in.  Most of the other pieces are in house with

the outstanding items being the glass cover and the windows for the vacuum

chamber.  These are scheduled to arrive next week.

A quick schedule was thrashed out in order to obtain an idea of the

timeline.  Assuming that we proceed with super-polished substrates of the

same size as those on the borrowed GEO output modecleaner, then we ought to

be able to deliver the hardware to the sites by the end of the first week

in September.  This also assumes that we can schedule a coating run right

after procurement of the substrates.

Optics Analysis

Erika D’Ambrosio

In collaboration with a few collegues in Livingston, we have worked 

on the shape of the beam that resonates in the power recycling 

cavity, when  the length of the resonator is smoothly changed.

There are two qualitative features that have been observed, which

were predicted by my theoretical analysis.

Basically everything can resonate in that cavity and, because of 

the geometrical coupling with the input beam, the power built up 

is not maximum if that coupling is not optimized, but the resonant 

curve is quite broad and the cavity is locked on a wide range of 

values of the common length. We will compare the optical gain

in a variety of situations and see how the sensitivity depends 

on the behaviour we are currently analysing.

Optical Contamination Cavities

Lee Cardenas, Liyuan Zhang

OTF Lab. (W. Bridge) 

The (21) pieces of the new twisted, shielded, teflon cables are under test.

We are taking ring down and beat frequency measurements 

as well as the RGA of the chamber every day. 

We have not noticed any change in terms of absorption nor contamination.

Absorption Test Measurement prototype in progress 

The sapphire substrate PINK color from Lyon, France. (314 mm Dia. X 131mm thick.) 

still inside the enclosure and it is in standby for further measurements.

We finished scanning the borrowed a small sapphire substrate rod 10mm in Dia.x 20mm long from Standford.

(~ 50-55 ppm/cm in loss at 1064nm.) 

I just finished making the support for the 80mm sapphire cube.We will calibrate this cube with the Sapphire rod.

We continue improving the absorption and detection sensitivity by testing different samples.

The fabrication of the rods for the enclosure is in progress.

Scatterometer system in standby

OTF Lab at Lauritsen ROOM 38 NO CHANGE.

Cavity #3 

The cavity is locked and we are taking ring down and beat frequency measurements every

day as well as the RGA for the chamber.

Cavity #2 Test cavity optical set up in standby ( NPRO s/n 414 was returned from 40m lab.)

We'll mounted back this NPRO and re-align it.

PSL LAB.

The 10 watt ND:YAG laser still in standby.

Misc... tasks 

The design and fabrication of thenew mount for the 80mm cube sapphire has been completed.

Spacer to increase the height of the mirror holder at the imaging lab is in progress.


40 Meter Interferometer (Weinstein)




Thermal Noise Interferometer (Libbrecht)


no report


LASTI (Ottaway)


no report


Data Analysis and Computing (Lazzarini)


Simulation and Modeling (Bhawal)
Weekly Physics meeting
--------------------------
We discussed sophisticated noise plotting into the existing LinLIGO
transfer
function model developed by Matt and the resolution problem for focussed
beam in FFT calculations.
 

Dual recycling field calculation
---------------------------------------
(Hiro) Still major effort was to complete this. The discrepancy between
 e2e primitive calculation and the matlab code using this fast summation
 is consistent with the accuracy of the approximation. Matlab code now
 has modulator / demodulator, and all necessary tools for simulation on
 matlab is ready.
 

 A LIGO note based on a mathematica analysis, T040062-00 has been
completed.
 This note, together with matlab code, e2e box files and some test
 results, will be available in the e2e home page.
 

Resolution for focussed beam
---------------------------------
(Biplab) Tried some adaptive method for increasing the beam resolution
for
 focussing beam in FFT way of propagation.
 

Code development and maintenance
----------------------------------
(Hiro)
 In order to support fast higher precision (higher than double) to cover
 wider dynamic range in the simulation, software tools were explored.
Two tools were available:  doubledouble
 (http://members.lycos.co.uk/keithmbriggs/doubledouble.html) by Briggs
 cannot be compiled by the latest gcc (was developed by gcc 2.7.2.1, and
 the author acknowledged this issue). The other was, qd
 (http://crd.lbl.gov/~dhbailey/mpdist/index.html) by LBNL group, worked
 but generates a slow code compared the native quad precision on intel
 linux box. We need to revisit this issue again to find a better
solution.
 

 This is an issue discussed at the March simulation meeting, and this
 was informed to Andrea Vicere of Virgo who is going to summarize the
 minute of the meeting.
 

(Melody) Modeler:
 Modifying the build scripts to change some default options and provide
 more user feedback.
 

 Continuing with the technical documentation for FUNC_X.
 

Alfi
-----------
(Bruce)
        - Redesign and implementation of connection and
          parser objects for use with new bundle/bundler objects.
 

LIGO Data Analysis System
 

Software Systems (Blackburn)
We have now rebuilt all of LDCG and LDAS using the GCC 3.3.3
compiler. Previously, only the C++ componentes of LDCG and LDAS
were compiled with the GCC 3.3.3 compiler. We are now testing
LDAS against this rebuild. All loop-script tests are running
as expected but we are still in the midst of performing the
broader system level tests. There are some indications that
LDAS is running a bit faster after this rebuild.
 

We completed testing of the timecheck flag in the createRDS
command and added a new check option to this command to allow
verification that the dataValid flag found in the FrADC data
channels are non-zero. Users may set to flag to allow non-
zero values, but the default is to through an exception if the
dataValid metadata is non-zero in the input data. This next
release of LDAS will have three new flags for quality assurance
of the input data as part of the createRDS command:
{-framechecksum, -frametimecheck, and -framedatavalid}.
 

We created a webpage that automatically captures and presents
the results of the nightly builds of LDAS. This webpage will
be linked off the main LDAS homepages shortly.
 

The standalone dataConditionAPI library is nearing readiness
for this release. The code will be alpha in nature, but we
expect to have a demo program to go along with distribution
as part of the LDAS 1.1.0 release.
 

We reduced minimum overhead associated with running a call
chain in the datacond API by 900 milliseconds. This was a
beneficial side effect of the code mods for logging the long
running thread, and it seems to have resulted in measureably
increased system performance.
 

The new spam filtering software installed by general computing
was complaining about the spamicity rating of the LDAS user
account email notification system. We made modifications to
this software to lower the rating below the current threshold
of 5 (prior to the changes the rating was 5.5). In addition,
similar changes were made to lower the emails from running
LDAS system from their previous level of 2.6 down to a level
of 0.5.
 

Added functionality to the database statistics plots in the
controlMonitorAPI to allow filtering results on particular
types of LDAS jobs.
 

Discovered that the major source of the long standing memory
leaks in the dataConditionAPI are a result of using the
database options {dbspectrum, dbqualitychannel, dbntuples,
dbquery}. This result came too late to allow fixing the
code for the 1.1.0 release, but it is great to finally have
a path to follow in tracking this problem. NOTE: some of
these commands have been seen to leak ~100KB/job.
 

We have now run the loop-scripts on a tandem system based
on the new DB2 8.1 database server. No unexpected issues
have been identified and we have now passed on the details
of using DB2 to Igor (at his request) for him to reopen the
effort of establishing a database federation/replication
solution for the project.
 

Worked on resolving about a dozen problem reports this week
of which about half were closed out.
 

The yearly LIGO-Griphyn face-to-face meeting took place at
ISI last week. The focus was on preparation and scheduling
for this year's supercomputing conference. The short summary
of the plan is to continue the effort on the periodic sources
and bring onboard inspiral searches for this year. A strong
effort will be made to handoff to the lsc community many of
the tasks that in the past required expertise from the ISI
computer scientist.
 

Hardware Systems (Anderson)
Caltech
-------
(Dan Kozak)
* Continued to try to track down SAM-QFS staging problem
  (staging stops working).
* Fixed problem with SAM-QFS taking drives offline.
* Began testing latest SAM-QFS patch using samtest with 9940A
  drives and T3 disks.
* Identified corrupt file in archive that seems to have been
  damaged during a grid-ftp transfer from LLO.  Replaced it
  with good copy from LLO.
* Worked with Sun to replace failed CPU in samtest.
* Continued deleting files in HPSS.
 

(Hari Pulapaka)
* Finished with setting up LDR data find server on ldas-gridmon.
* Adding new entries in the rls server for the LDR data find.
* Experimenting with psyco, a module for increasing python's
  speed (JIT).
* Helping out Keith and Igor with VDT setup at the sites.
 

(Al Wilson)
* Switching the datacahes to use the command line interface for
  media checking rather then web based tool.
* Setting up BB to read the output from the CLI.
* Clearing Problem reports.
* Ldasbox1 - Ed reported the box1,2 both are running at 800MHz
  not at 2.4GHz, still looking into this problem.
 

(Stuart Anderson)
* Renewed the Sun T3 contract for 53 units.
* Renewed the Foundry contract for the large copper GigE switches.
* Continuing to interleave the S2 and S3 level-1 RDS frames
  across the internal node disks in the LDAS-CIT cluster.
* Started testing Linux kernel 2.6.5 for NFS server performance.
* Installed new 19" equipment rack in 605 Millikan.
 

MIT
---
(Keith Bayer)
* 4 more 20amp power circuits being brought into ldas lab.
* Working on VDT install on cluster nodes.
* Working on Condor install on cluster nodes.
* Working on resetting LDR database with current datasets at MIT.
 

Hanford
-------
(Ben Johnson)
* Worked with Al Wilson to set up routine media scans on the
  IDE RAID units at LHO. I'll write up some cronjobs that will
  run the scans this Sunday.
* Replaced a failed disk in the new IDE RAID here. This was
  done in a cold swap fashion, but some command line features
  were tested. They appear to work as expected.
* A summary of findings for the data corruption problem was posted
  in the LHO e-log this past Friday, April 2. Continuing work on
  final analysis of the problem in addition to the final report.
 

Data Analysis Activities (Lazzarini)
 

Creighton:
This past week I attended the GriPhyN/LIGO face-to-face at ISI, and
have been poking at an analysis of signals from thermal-instability
r-mode bursts from recycled neutron stars.
 

Shawhan:
* Finished paper for GWDAW proceedings about a new waveform consistency
test for inspiral searches (work done with Evan Ochsner [SURF student]
last summer)
 

Sutton;
1. Kicking-off the code review for WaveBurst.  (This is the first
review of
the untriggered burst codes.)   Michele Zanolin and I are conducting the
review.   We began with a telecon with the Burst review committee (Riles
et al) and with Klimenko et al to plan the review.
2. Continuing collaboration with Yakushin to organize the production of
MDC
frames for S2 LIGO and LIGO-TAMA analyses.
3a. Drafted and circulated an article on the plans for LIGO-TAMA for
submission to the GWDAW proceedings.
3b. Also doing Monte Carlo simulations of the LIGO-TAMA network
efficiencies as a function of the efficiencies of the individual LIGO &
TAMA
IFOs.  This is intended as guidance for tuning of the LIGO-TAMA
analysis.
 

Sylvestre:
o I have completed running over the S2 injection frames for
sine-gaussians
using TFClusters. The frames had a (known) problem, but this shows that
the method works. It takes ~3 days to produce a list of events that can
be
used by the r-Statistics.
 

o I have been doing some preliminary work on a new scheme to extract
waveform information from our time series. This might be useful for
consistency checks on events passing cross-correlation tests, for
instance.
 

o I have also been cleaning up some of my codes, in preparation for the
code review of the TFClusters ETG.
 

Weinstein:
- Working on issues related to inspiral analysis review.
 

Yakushin;
1) Debugging the new MDC frames script for LIGO and LIGO-TAMA analysis.
2) Together with S. Klimenko and M. Rakhmanov preparing GDDAW-8
presentation on waveburst for publication in Classical and Quantum
Gravity.
3) Helping Michele Zanolin to learn to use waveburst for the internal
burst group software review.
 

Lazzarini:
Working on time-dependent stochastic background detection.
 

General Computing (Wallace)
MIT:
(Keith)
-Experimenting with AIT3 compression for tape backups
-Investigated network loss in NW22 (circuit breaker tripped
in plasma student office that also affects power to our
fiber link transceiver in adjacent router closet.
 

Livingston:
(Shannon)
-Current bandwidth usage can be seen at
http://teche.ligo-la.caltech.edu/mrtg/LLO-Router/130.39.245.1_1.html
-Archived bandwidth usage can be seen at
http://teche.ligo-la.caltech.edu/mrtg/archive/ under the dated folder
for the week of interest.
-Moved the mail server, file server, support web server, and another
machine to a new rack in the computer users room.  Also moved the Gb
switch into the new rack.  There is  a larger UPS in the rack also
which should run the servers for ~20-40 minutes instead of ~5 for the
old one.
-Setup a new laptop for Bonnie.
-Moved all of the shared application installation files from a CD
server to our main file server.
-Made further firewall changes for LDAS.  This is to support the new
servers LDAS has installed at LLO.
-Spent a considerable amount of time installing the new rack and
rearranging the computer users room to accommodate the new rack and UPS
in the computer users room.  This has been needed for a long time.
Finally I can move all of the servers out of the Comm room and into a
proper rack.  There is one rack in the room now, later this week a
second rack will be available.  Both racks will now have better UPSs
and network switches.  I will have to order a new KVM however.
-Working on providing more network ports to the new building.  I have
finally run out of switch ports there.  This is causing a problem since
I need a couple more wall jacks activated.  I am working on  a short
term solution since delivery on a fiber I need is several weeks out.
-Setting up a couple of rack mount computers for the Intellution
software.  These will be shipped to the contractor and returned with
the necessary software installed.
 

Hanford:
(Christine)
- Network bandwidth usage can be seen at
http://apex.ligo-wa.caltech.edu/~christin/mrtg/ 198.129.208.1_198.129.78.122.html
- Completed installing software on the license server.  I'll be ready to
swap over by the end of the week.
- Installed Win2k on a laptop and on a dell server for CDS.  Helped with
network problems on the laptop and other software problems on the
server.
- Spoke with a PNNL representative about the change in network
connection from OC3 to GigE.  Looks like PNNL/ESnet are planning to give
us a GigE connection at PNNL and wanted to make sure I had the correct
adapter card for my router.
- No word on when our extra pair of fibers will be connected at PNNL.
The phone company who will be doing the work is busy with a major
hardware swap out at their office.
- Fixed some network and printer problems on one of the community Sun
computers.
- Misc. user support.
 

CIT:
(Mike)
-Loaded Solid Works on Mike Smith's workstation. This took a lot of
time,
due to incompatibility issues. I found some updated drivers for the
problem
video card, that worked.
-Finished up loading a new laptop for Peter Shawhan. I loaded all GC
software and setup network and security settings.
-Worked on the color printer, located up on Millikan. This printer had
many
problems with transfer belt and power cycle errors. I have this printer
back up and running. The black & white printer also had many problems
that
I corrected.
-I worked on a laptop for Magdalena & Barbara Simoni. They had many
problems with Adobe & Microsoft office. After making many adjustments, I
finally figured out how to get this software to work again.
-I swapped out a color (hp4500n) printer that is located on the 2ed of
W/B.
I swapped this out with a spare 4500 printer. This also included me
having
to swap hardware from the old printer, in order to get the spare printer
working. This is now back up and running.
-Worked on Fumiko Kawazoe's workstation. This required a rebuild plus
hardware upgrade. This computer would not stay booted, it just continued
rebooting itself. I upgraded some of the hardware, with parts from a
surplus computer and transferred over users data. She is back in
business.
-Loaded a new workstation for Calum Torrie. This reload is very time
consuming, due to the multiple engineering packages that are required
for
this type of workstation. I then transferred all this users data over,
from
old workstation. I also swapped out his old setup/workstation and setup
the
new workstation in his cubicle.
-Installed a FlexLm Server for Algor. This license server is running on
(pictor)
-I worked on the NTSVR's updating critical updates and then looking into
event viewer for system errors. I also made some backup changes on M91
the
media server. I coordinated this with Veronica/Larry.
-Worked on other misc. user support.
 

(Veronica)
- LSC website: Started working on a website for the upcoming LSC June
meeting. Looking into ways to integrate an ecommerce application system
for online payments set up at MIT into the website.
- CaJAGWR website: Updates to the seminars page. A follow-up on the
video
of the last CaJAGWR/PCGM seminar; the video was shot in High8 format. I
had it converted to miniDV. I then recaptured and compressed it and
posted
a pointer to the video stream.
- LIGO website: Did a massive amount of media backups for files older
than
one year, which were captured directly onto hard disks and for which no
tape footage exists that could be used as a backup. Updates to various
pages and usual upkeep.
 

(Lisa)
- Updated the canit software on both mailserver. I have been spending
2-4 hours
of every day, this past week, customizing the filters and looking for
false
positives.
-Spent some time working on my tcl script to program the AP-2000.  I
need to
get this script in place before finishing the rollout or maintenance
will be a
real problem.
- Spent more time working out the glitches with the mailman server
following the
mailserver work.
 

(Larry)
-We have had a number of requests for new accounts. There are people
from the
LSC as well as students for this years SURF program all needing
accounts. So,
far no major problems but as a reminder everyone not directly part of
LIGO needs
to have a sponsor and everyone needs to signoff on the LIGO GC computer
policy,
before receiving and account.
We have started working on some of the logistics to try and make sure
we have
computers for the incoming SURF students and visitors.
-Went through the E2E computers again. One unit is going to need a
rebuild. It
is running but doesn't appear to be too stable. All of the units are
having
timing issues so some sort of NTP service will be setup on the systems.
The decision has been made to have a separate application server as
part of the
server cluster.
-Mathematica has come up with a font problem. Running it on a local SUN
box has
no issues but trying to run it remotely does. The exceed s/w has a
known problem
and Mathematica has some fonts that need to be installed with the
exceed s/w.
Albert has put in a trouble call for a similar problem when trying to
run
remotely using ssh. We've run a number of tests and found that if you
have run
mathematica locally and then remotely, you do get some error messages
but things
do run the way they are suppose to. We will continue working on the
problem.
-Worked a little with some of the problems with IDEAS. It appears that
Caltech
no longer has a site license for the pkg., a call has been put in to
see what we
can do about getting the application license. Dennis C. has been
working a
number of other issues to get his analysis to work properly with the
pkg.
-Worked a number of orders for different people.
Assisted Lisa and Cleveland in shipping back a couple for computers
used for a
trade-in purchase.
-Assisted a number of users with different issues. Most of the problems
turned
out to be typos.
-Still working on the computer room expansion. PMA is working on
getting in some
quotes for the work. Did a little more work on the racks in the new
area if the
computer room.
 


Advanced LIGO Development (Shoemaker)



 

Advanced LIGO and supporting R&D

Seismic Isolation

From: Larry Jones <ljones@ligo.caltech.edu>

Advanced LIGO Seismic Structure

SEI Structure Design/Fab contract:

We have decided to delay the fabrication of the HAM SEI structure prototype, in order to minimize the cost growth of this contract. HAM design work will be completed, to capitalize on the excellent work being done in analysis and design and to permit a smooth transition into fabrication when funds permit.

Work continued on detailed finite element model of the BSC support structure. Coyne and Jones joined Tony Agajanian of ASI for a tour of Astro Pak, ASI's cleaning vendor.

The Design Review meeting for the BSC structure is now scheduled for May 18, 2004. The HAM structure Design Review is scheduled for July 30, 2004. Delivery at LASTI of the BSC prototype is expected to be in November, 2004.
 

Actuator Testing and Redesign:

The Change Board approved procurement of the actuators, and a revised quote was received for the reduced quantity (for the BSC structure prototype only). The contract is being prepared, which will include procurement of design drawings for competing the production quantities later.
 

Position Sensor Probe Development:

The Change Board approved procurement of the position sensors and related equipment; details are being finalized with the vendor, and a reduced quantity quotation has been requested.
 

Seismometer Procurement:

Information is being collected on Streckeisen's plan to replace the STS-2 model with the new STS-3, in an attempt to minimize the difficulty of production modifications. The prototypes will use STS-2 units only.
 

Galling/Dusting Test:

All testing is now complete. Reporting is awaiting time available.

From: ctorrie <ctorrie@ligo.caltech.edu>

Weekly Report
Advanced LIGO Suspensions

1) FEA
Work is continuing to support Dennis on the structure and table analysis.

2) Controls Prototype Quad
A draft of the requirements of the final stages of the two chains for the quad have been created for the controls prototype and will be distributed once I have updated with comments HA, JHR and NAR.
Using SolidWorks Explorer I have created an assembly for each of the four masses: penultimate mass, test mass, reaction mass and penultimate reaction mass. this work is ongoing.

3) Prototype Single and Double Pendulum
Parts are now in workshop to expand the single pendulum experiment to include the test of a 4x4 array eddy current damper on a 40kg mass.

4) Prototype Catecher Experiments
Parts are in workshop to cover a series of experiments laid out between CAC, JHR, RJ and myself to aid in translating from a GEO style catcher to one for the LASTI quads.

Suspension

From: JaneenRomie <romie_j@ligo.caltech.edu>

AdLIGO Suspensions
Working with Dennis and Calum on Quad/SEI fea.

Made mock-ups of two upper masses and two blades for the quad.

Participated in the osem conceptual design review.Working on actions from the meeting.

Sent out a request for quote for a suspensions gazebo to a vendor in New York. They will quote for an assembly/kit made up of Bosch-like aluminum extrusions. We will add a .25" thick x 4' x 8' alum. plate as a dummy BSC optical table.

Had a discussion with Russell, Caroline and Calum about catcher design.

Have received some comments on the Suspensions Controls Prototype Test Plan and will start to incorporate additions.

Worked with Thomas this afternoon about incorporating Primavera tracking info on the quad task list. I will send a first cut to Thomas at the end of this week. Primavera updates will be done after that.


Pre-stabilized Laser

From: Peter King <pking@ligo.caltech.edu>

PSL
===
    No apparent problems have been observed with the NPRO laser that was
removed from the 40-m Lab 10-W laser.  When inspected at random intervals
the output power was within 2 mW of its setpoint and the crystal
temperature was within 0.2 degC of its setpointHowever when the crystal
temperature and current, and the diode temperature and current signals are
displayed on an oscilloscope there is a strong ~20 MHz signal present.

This masks any glitching in the signals, if there is any so I have not been
able to compare these signals with those posted in the LLO e-log.

    A broken capacitor --- 10 pF I think --- was found on the current shunt
board.  This was replaced and made a small difference to the transfer
function around 2 MHz.

OMC
===
The piece for the output modecleaner vacuum chamber was EDMed okay.
Part of the delay was in finding a local wire cutting shop that could
handle pieces larger than 6 in.  Most of the other pieces are in house with
the outstanding items being the glass cover and the windows for the vacuum
chamber.  These are scheduled to arrive next week.

A quick schedule was thrashed out in order to obtain an idea of the
timeline.  Assuming that we proceed with super-polished substrates of the
same size as those on the borrowed GEO output modecleaner,then we ought to
be able to deliver the hardware to the sites by the end of the first week
in September.  This also assumes that we can schedule a coating run right
after procurement of the substrates.

AdvLIGO PSL
===========
    I met with Thomas last Friday and handed over some revisions to the
cost book.  Incorporated in the latest revision is the proposed split along
the lines of who pays for what, as this picture changed from previous
notions.  Under the latest revision, LIGO Lab will be responsible for a
greater share in the PSL cost.


Core Optics

From: Helena Armandula <ahelena@ligo.caltech.edu>

Adv.LIGO Coatings
J.M.Mackowski confirmed his trip to the States. He'll be visiting Caltech during the week of April 19th; the date has yet to be confirmed. A teleconference will be setup to take advantage of his visit to discuss coating issues.

AdvLIGO SUS
Researching methods that would allow to silicate bond large pieces of glass minimizing defects such as bubbles. Moderate progress achieved.

From: Bill Kells <kells@ligo.caltech.edu>

I am planning some "last ditch"
measurements/investigations to try to
arrive at further insights into the FS vs Sapphire
choice. This will involve closely looking at the
bulk inhomogeneities of samples we have now with
the instruments we have operational in our OTF. Will
be able to say more next Thrsday.




Other Laboratory R&D


From: RiccardoDeSalvo desalvo@ligo.caltech.edu

I resume the SAS group weekly after several weeks.I have been travelling and/or too busy for it.Maddalena has collected weekly entries from the other people of the group and fed them to the LIGO weekly, but not to the usual mailing list.Unbelievably I got complaints!I apologize.

Before the reports from the rest of the group, I summarized my activities during and after the travel.Some of this may have already reached you through Maddalena's reports.I made lots of presentations, some are fun to watch, some I may have reported already.I also asked the rest of the group to summarize their work for the past few weeks.

In Aspen I presented some of the group results,Bodies in Motion, about Maddalena’s new springs, available at 

http://www.ligo.caltech.edu/~desalvo/presentationLIGO-G040029-feb-04/

including movies,

and “mining for Gravitational waves” on low frequency GW detection

available at:

http://www.ligo.caltech.edu/~desalvo/G-040036-00-R.ppt

The results from this talk may open the way to underground, LF (1 to 20 Hz) GW detection.

In China, together with Ren-yuan Zhu I participated to the Kickoff workshop of CEGO, the Chinese, Einstein, underground Gravitational Wave Observatory.They are interested in building their own underground GW detection facility.Incredibly enough we did not suggest them, it was just pure serendipity!

In ChinaRen-yuan Zhu presented the status of advance of LIGO, his presentation is available at:

http://www.ligo.caltech.edu/~desalvo/ryz_ligo-G040222-00-R.pdf

I presented a talk on challenges and rewards from Gravitational Wave

LIGO-G040049-00-R available at:

http://www.ligo.caltech.edu/~desalvo/MARCO-POLO.ppt

As the Chinese are not expert in the field, they asked for our support.LIGO and other GW experiments, as well as GWIC responded positively, contacts are ongoing with LIGO, NSF and other institutions to decide on the level of the possible external support that can be provided.

Yanbei Chen and myself later contributed editing the technical base for the CEGO proposal to the Chinese government see draft at the bottom.

In Japan I visited Ryutaro Takahashi at TAMA.He received the sensors, actuators and electronics for the TAMA-SAS, the mechanics will arrive mid April, then SAS assembly will start.I also gave the two Aspen talks in TAMA and Hongo, as well as a presentation of the mexican hat mirror experiment,LIGO-G040223-00-R available at:

http://www.ligo.caltech.edu/~desalvo/Mex-Hat-presentation.ppt

At LSC I presented the Chinese project and their request for support,the talk is available at

http://www.ligo.caltech.edu/~desalvo/PresentationLSC3-2004.ppt

At Granlibakken, at the seismometer R&D workshop, I got informations on flex joints and gave a short presentation of LIGO, mostly unknown in the gophysicist community.

http://www.ligo.caltech.edu/~desalvo/granlibachen.ppt

When (briefly) back in Campus, Yanbei Chen and myself edited, with the help of many, a draft of a proposal which will be, translated in Mandarin and reworked, the base of the proposal to the Chinese government for the initial request of funds, for the studies, R&D and experiments, in view of the design and final proposal of the real facility.

It was a heavy job, the draft is available 

http://www.ligo.caltech.edu/~desalvo/proposal_yanbei-ric-4.doc

I am happy it is done.

In view of veryfying the assumptions in “Mining for Gravitational Waves” FabrizioBarone, GC. Cella and myself are organizing a measurement campaign in July in the Realmonte mine access tunnel to measure the coherence and strength of seismic waves as a function of depth.

The activities on the flat beam profile interferometers are listed below.

About the two papers ready for publication, Hareem and Virginio are waiting for the reply from the internal referees.

Juri, Erika

I am going on with the study of Mexican Hat mirrors cavities and the feasibility of a dynamical characterization (using some tools for mechanical simulations) of the rigid cavity that is going to be built. In particular I had a useful discussion with Prof Yamamoto who gave me some issues to think about and now I have some constraints we must take into account. I spent also some time with the Mathematica's files PavlinSavov gave me; these are used to find the eigenmodes of a Mesa Beam in the case of flat MH mirrors or for MH concentric cavity; I made some improvements for the numerical calculations.

I obtained the shape of the flat-topped beam inside a concentric cavity using Mathematica's calculation tools in which I used dimensionless length-units, scaled over the wavelength of the laser, in order to improve the numerical calculations.Than I found the mirror's surface correction in order to support this mesa-beam in the case of the small cavity that we building. Now I'm going on with the calculations to obtain the eigenvalues and eigenvectors of this Concentric-Mexican-Hat cavity that will be useful to study the stability of this configuration. I asked EnricoCampagna some information about Bench that I will use in the future to calculate interferometer performances.

I worked with Erika D'Ambrosio to find an analytic proof of a kind of equivalence between flat and concentric cavity that was pointed out also by our numerical calculations. We found a very elegant relation between the kernels,eigenvalues and eigenvectors of the two cavities that can be extend also to the mesa-beam configuration of both cavities. In this way we can design the mirror shape of the MH-concentric cavity, find its eigenvalues and eigenfunctions, simply using the calculations done for the flat case.The key point is that for a confocal cavity the mexicanhet corrections are negative instead of positive.While a flat mexican hat mirror has a rim sticking out, a spherical mexican hat mirror has a bevel.Now I am calculating the best spherical mirror necessary to fabricate the 7 meter interferometer mirrors by corrective coating in order to generate a MH-concentric cavity (obviously it would be very difficult to grind an accurate negative correction to the mirror's surfaces). We need to be ready in a week from now because J.M. Mackowsky is stopping by to finalize the design.

We are also writing a report about our study of the equivalence between MH-flat and MH-concentric cavities and the angular instability problem.

Maddalena

I have analyzed the Q factor behavior of MGAS blades by changing the mechanical stiffness of the filter by changing the radial compression of the blades in two or three positiona and then by producing the Q-to-frequency curves by changing the stiffness electrically using the control circuit. The behavior of the Q factor in function of the resonant frequency has the same power value for different mechanical stiffness, the only change is in a multiplicative factor.

Continued to make filter passive attenuation transfert functions. In this set of transfer function I have obtained resonant frequencies below 80 mHz, but I could not go at lower frequency because the filter working point depends strongly, at these low frequency, on the temperature changes and goes easily out of the working point. 

To avoid the thermal problem I’m now making a thermal compensation circuit to maintain the spring at the working point.

After that we will improve the mechanical setup to reduce the asymmetry and balancing the system with a coaxial system of LVDT and actuator. 

For this reason I’m calibrating the new actuator and making some modifications to better linearize its responce.Since we are using the TAMA-SAS actuators, the modifications can also be applied to the TAMA, Napoli and Firenze TAMA SAS towers.

Barbara

In the last few weeks, after having tested the behavior of the test maraging springs, suspending a weight equal to 1/4 of the cavity and finding the working point for two different stiffness values of the blade.An error in the initial blade’s design forced us to enlarge the vacuum bells housing the cavity suspensions.Happily the bells were not made yet.We finally approved the final design for the mechanics of the cavity implementing also Maddalena’s low frequency control and thermal compensation of the suspension system adding an LDVT, an actuator and a tuneable correction spring on each suspension.

After the approval of the design and making the final draft we ordered the mechanical parts and components including the stepping motors to drive the suspension system correction spring and the connectors to drive the piezoelettric and all the components inside the vacuum.

We are clearing up the OTF lab to host the 7 meter cavity.

For what concern the mechanical parts, this week we have received the first part of the mechanics (Invar rods and rigid cavity structure, piezos…), and in the next 2 week we will get ready to begin to build the cavity itself.The suspensions and the modified vacuum, delayed by our mistake, are due in a couple of weeks.

At the same time we expect to receive the spherical optics to begin to work with our cavity.

About the final configuration of the mexican hat cavity, few days ago we investigated the possibility to use a confocal cavity instead a flat one (to match the new Advanced LIGO design), and we have decidedthat we will probably begin with the flat mirror (to be cross checked with J.M. Mackowsky), after which we would switch to the spherical mexican mirrors to analyze both configurations.

Also in the last few weeks I wrote some slides to show the status of the apparatus and parts of my thesis.

 

 


For additional information about this report, contact sanders@ligo.caltech.edu